# Ytterbium atom interferometry for dark matter searches

### v1

@@ Abstract
We analyze the sensitivity of a laboratory-scale ytterbium atom interferometer to scalar, vector, and axion dark matter signals. A frequency ratio measurement between two transitions in 171Yb enables a search for variations of the fine-structure constant that could surpass existing limits by a factor of 100 in the mass range $10^{-22}$ eV to $10^{-16}$ eV. Differential accelerometry between Yb isotopes yields projected sensitivities to scalar and vector dark matter couplings that are stronger than the limits set by the MICROSCOPE equivalence principle test, and an analogous measurement in the MAGIS-100 long-baseline interferometer would be more sensitive than previous bounds by factors of 10 or more. A search for anomalous spin torque in MAGIS-100 is projected to reach similar sensitivity to atomic magnetometry experiments. We discuss strategies for mitigating the main systematic effects in each measurement. These results indicate that improved dark matter searches with Yb atom interferometry are technically feasible.
@@
Phenomenologically, the signal in a dark matter direct detection experiment changes qualitatively at a dark matter particle mass of $\sim 1$ eV. If the dark matter particle mass is above $1$ eV, the phase space density of the dark matter near the Earth is less than one, and the dark matter is expected to exhibit particle-like behavior (e.g., scattering off of ordinary matter). Several experiments Alkhatib et al. (2021); Abdallah et al. (2015) have searched for dark matter in this regime. If the dark matter particle mass is below $1$ eV, however, the phase space density of the dark matter near the Earth is greater than one. In this case, the dark matter must be a boson, and its behavior is analogous to that of a classical field: all of its lowest-order couplings to the Standard Model yield signals that oscillate at the Compton frequency of the dark matter Graham et al. (2016a); Arvanitaki et al. (2015). Experiments searching for this “ultralight” dark matter include optical clocks Filzinger et al. (2023); Sherrill et al. (2023), microwave cavity experiments Bartram et al. (2021), nuclear magnetic resonance experiments Aybas et al. (2021); Wu et al. (2019), and atomic magnetometers Bloch et al. (2020); Lee et al. (2023), among others. Ultralight dark matter would also give rise to new static forces between ordinary matter. Such forces have been constrained by equivalence principle tests Touboul et al. (2022); Schlamminger et al. (2008).
@@
In this work, we consider the potential physics reach of dark matter detection experiments based on atom interferometry. In a light-pulse atom interferometer Hogan et al. (2009), ultracold atoms are split by atom-light interactions into a superposition of external states, which are then recombined and interfered. Depending on the interferometer geometry, the phase of an atom interferometer can be sensitive to inertial forces Kasevich and Chu (1991), recoil velocity Bordé (1989), and/or the energy difference between internal states. Atom interferometers have been used to test the equivalence principle at a relative accuracy of about $10^{-12}$ Asenbaum et al. (2020), measure the fine-structure constant Morel et al. (2020); Parker et al. (2018), and observe gravitational phase shifts in nonlocal quantum systems Overstreet et al. (2022); Asenbaum et al. (2017). In addition, several gravitational wave detectors based on atom interferometry Yijun Jiang et al. (2021); Badurina et al. (2020); Canuel et al. (2018) are under construction.
@@
High-precision atom interferometers typically have a coherence time on the order of $1$ s and a cycle time of a few seconds; thus, atom interferometers are naturally sensitive to dark matter with Compton frequency $\lesssim 1$ Hz. With a measurement campaign of about one year, an atom interferometry experiment can search for dark matter signals over eight orders of magnitude in the dark matter particle mass (from $10^{-8}$ Hz to $1$ Hz, or $10^{-22}$ eV to $10^{-14}$ eV). The possible signals from ultralight dark matter depend on its spin and parity. A spin-zero or spin-one dark matter field can produce six qualitatively different signals at lowest order Graham et al. (2016a). Of these, atom interferometers can be sensitive to three: variations of fundamental constants, accelerations, and spin torques.
@@
One of the main considerations for an atom-interferometric dark matter detection experiment is the choice of atomic species. Ideally, the experiment should use an atom that provides high sensitivity to dark matter signals while minimizing systematic effects. Here we consider atom interferometry with ytterbium isotopes. Ytterbium offers the highest sensitivity of any neutral atom to variations of the fine-structure constant in one of its excited states Safronova et al. (2018); Dzuba et al. (2018), possesses multiple isotopes that can be cooled simultaneously for a differential acceleration measurement Kitagawa et al. (2008), and has an isotope with nuclear spin $1/2$ that can be used to detect spin torque. In addition, the alkaline-earth-like electronic structure of ytterbium provides magnetic insensitivity in the ${}^{1}S_{0}$ electronic ground state, and its multiple narrow-linewidth transitions facilitate precise measurements of transition frequencies.
@@
The remainder of this paper is organized as follows. The apparatus is described in Section II, and its sensitivities to scalar, vector, and axion dark matter are discussed in Sections III.A, III.B, and III.C, respectively. Section IV describes the main systematic effects in these measurements and how they will be controlled. Section V compares atom-interferometric dark matter searches to other experiments and discusses the prospects for future sensitivity improvements.
@@
The JHU experimental apparatus will consist of a source of ultracold Yb and an atomic fountain in which the atoms can freely fall during interferometry sequences. Clouds of ultracold Yb will be produced in a 3D magneto-optical trap (MOT) that is loaded by a commercial Zeeman slower/2D MOT. The 3D MOT will utilize core-shell techniques Lee et al. (2015) to increase atom number and phase space density. The atoms will be evaporatively cooled in an optical dipole trap and then launched into a magnetically shielded atomic fountain by an optical lattice. The height of the magnetically shielded region will be $2.5$ m, enabling interferometry over a free-fall distance of $2$ m. After the interferometry sequence is complete, the interferometer phase will be measured via fluorescence detection of the number of atoms in each output port.
@@
For the purpose of generating estimated sensitivities to dark matter, we assume that $10^{5}$ atoms will participate in each measurement. We also assume that the coherence time of each measurement will be $1.3$ s, limited by the available free fall time, and that the experimental cycle time will be $10$ s. These atom numbers and cycle times have previously been achieved in high-precision atom interferometers Asenbaum et al. (2020).
@@
Atom interferometers naturally detect energy level shifts and accelerations. With the appropriate geometry, an atom interferometer can be sensitive to scalar, vector, or pseudoscalar (axion) couplings to ordinary matter. In this section, we describe interferometer geometries suitable for detecting each of these dark matter candidates and calculate the projected sensitivity of the JHU apparatus to the associated coupling. For each estimate, we assume shot-noise-limited statistics and a one-year measurement campaign. The leading systematic effects for each measurement are discussed in Section IV.
@@
$\alpha(t)=\alpha_{0}\left(1+d_{e}\varphi_{0}\cos\left(\frac{m_{\text{DM}}c^{2}}{\hbar}t\right)\right).$ (3)
@@
A frequency ratio measurement between two transitions is sensitive to variations of $\alpha$ as long as the value of $\Delta q$ differs between them, and transitions with large values of $\Delta q$ have the highest discovery potential.
@@
We propose to use a 171Yb atom interferometer to measure the frequency ratio between the ${}^{1}S_{0}\leftrightarrow{}^{3}P_{0}$ transition at $578$ nm and the ${}^{1}S_{0}\leftrightarrow 4\text{f}^{13}6\text{s}^{2}5\text{d}\,(J=2)$ transition at $431$ nm. The energy of the $4\text{f}^{13}6\text{s}^{2}5\text{d}\,(J=2)$ state has the highest sensitivity among neutral-atom states to variations of $\alpha$ Safronova et al. (2018). An interferometer geometry for this frequency comparison is depicted in Fig. 1. In each of the four Ramsey-Bordé interferometers, a single-photon transition creates a superposition of internal states, which are interfered to produce an interferometer phase of the form Hogan et al. (2009)
$\phi=2\omega T+\frac{\hbar\omega^{2}}{2mc^{2}}T+\frac{\omega}{c}gT^{2}+\cdots.$ (5)
Here $\omega$ is the transition frequency, $T$ is the interferometer time, $m$ is the atomic mass, and $g$ is the magnitude of the laboratory acceleration relative to a freely falling geodesic. The first term is the desired signal, while the second and third terms represent systematic effects due to the recoil velocity and the laboratory acceleration, respectively. To suppress these effects, we consider the differential phase between pairs of interferometers [(a) - (b) and (c) - (d)], which has the form
$\Delta\phi=4\omega T+\cdots$ (6)
and is insensitive to the phase shifts arising from the recoil velocity and the laboratory acceleration because they are common to both interferometers. An alternative geometry for measuring this frequency ratio is shown in Fig. 2, where each beamsplitter is implemented by a Doppler-free two-photon transition driven by counterpropagating laser beams. Compared to the geometry in Fig. 1, Doppler-free transitions naturally suppress the recoil shift, thereby reducing the number of interferometers required from four to two, and prevent atom loss into undesired momentum states. However, Doppler-free two-photon transitions require higher laser power than single-photon transitions for a given Rabi frequency.
@@
Here $r$ is the distance between particles, $m_{i}$ and $Q_{i}$ are the mass and the dilaton charge of particle $i$, respectively, and $\lambda$ is the Compton wavelength of the dark matter particle. Note that the coupling constant $d_{e}$ in this expression is the same quantity as in Eq. 1. Since the dilaton charge of an atom is a function of its mass and atomic number (see Appendix A.I), this potential leads to a composition-dependent static force
@@
where $n$ is the number of photon recoils in the initial beamsplitter, $k$ is the magnitude of the laser wavevector, and $\Delta g$ is the acceleration difference between isotopes projected onto the interferometer axis. For this measurement, we consider a comparison between 170Yb and 176Yb, which are both bosons and have favorable scattering lengths Kitagawa et al. (2008) that allow simultaneous evaporative cooling. The beamsplitters can be implemented by Bragg transitions on the electric dipole transition at $399$ nm or Bragg transitions on the intercombination transition at $556$ nm. We also assume $n=500$, allowing a maximum wave packet separation of $2$ m at time $T$. Beamsplitters with hundreds of photon recoils have previously been demonstrated Wilkason et al. (2022); Béguin et al. (2023). With these parameters, the JHU apparatus is projected to reach a sensitivity to the Eötvös parameter $\eta\equiv\Delta g/g=2\times 10^{-16}$ after a one-year measurement campaign.
@@
Figure 4: Constraint plot and projected sensitivities to scalar dark matter coupled to the electromagnetic field tensor. JHU $\alpha(t)$: projected sensitivity of spectroscopic search in JHU apparatus with interferometer geometry shown in Fig. 1. JHU (DC): projected sensitivity of static equivalence principle test in JHU apparatus with interferometer geometry shown in Fig. 3. MAGIS-100 (DC): projected sensitivity of static equivalence principle test with Yb isotopes in MAGIS-100 long-baseline interferometer. Also shown are existing constraints from the MICROSCOPE experiment Touboul et al. (2022) (purple curve) and Yb ion clock experiments at NPL Sherrill et al. (2023) (green curve) and PTB Filzinger et al. (2023) (yellow curve). Shaded regions are excluded by existing constraints. All curves are 95% confidence limits.
@@
Figure 5: Constraint plot and projected sensitivities to scalar dark matter coupled to the electron mass. JHU (DC): projected sensitivity of static equivalence principle test in JHU apparatus with interferometer geometry shown in Fig. 3. MAGIS-100 (DC): projected sensitivity of static equivalence principle test with Yb isotopes in MAGIS-100 long-baseline interferometer. Also shown are existing constraints from the MICROSCOPE experiment Touboul et al. (2022) (purple curve) and a hydrogen maser-optical cavity comparison Kennedy et al. (2020) (green curve). Shaded regions are excluded by existing constraints. All curves are 95% confidence limits.
@@
The projected sensitivities of these measurements to the scalar dark matter coupling constant $d_{e}$ is shown in Fig. 4 as a function of the dark matter particle mass. The projections are compared to existing constraints from Yb ion clock experiments Filzinger et al. (2023); Sherrill et al. (2023) and the MICROSCOPE space-based equivalence principle test Touboul et al. (2022). We calculate the projected sensitivity to variations of $\alpha$ both analytically and by taking the discrete Fourier transform of simulated data. In the JHU experiment, the Lomb-Scargle periodogram VanderPlas (2018) will be used instead of the discrete Fourier transform to optimally account for variations in cycle time and experimental dead time. The projections take into account the stochastic amplitude of the dark matter field, which diminishes sensitivity to oscillating signals by a factor of 3 Centers et al. (2021). At higher dark matter masses, the search for variations of the fine-structure constant loses sensitivity due to signal averaging over the $\sim 1$ s coherence time of each experimental run, while the equivalence principle test loses sensitivity due to the exponential decay term in Eq. 8, which effectively limits the quantity of material in the Earth that sources the static force (see Appendix A.II). We also plot the projected sensitivity of an equivalence principle test performed between 170Yb and 176Yb in the MAGIS-100 long-baseline interferometer at Fermilab Yijun Jiang et al. (2021). For this projection, we use the interferometer geometry in Fig. 3 and the same experimental parameters as in the JHU apparatus (atom number, photon recoils, etc.), except that the interferometer time is increased from $T=0.64$ s to $T=4.5$ s. We note that this measurement in MAGIS-100 would reach an Eötvös parameter sensitivity of $\eta=5\times 10^{-18}$.
@@
The JHU atom-interferometric search for variations of the fine-structure constant is projected to improve on existing experiments by two orders of magnitude in the mass range $10^{-22}$ eV to $10^{-16}$ eV. The Yb equivalence principle test performed in the JHU apparatus is projected to reach a similar sensitivity to the MICROSCOPE experiment, while the MAGIS-100 equivalence principle test would be a factor of 10 more sensitive than existing bounds from $10^{-16}$ eV to $10^{-12}$ eV.
@@
where $\hat{\mathbf{e}}$ points in the polarization direction of the dark matter field. In addition, atoms can exert a force on one another by exchanging the vector particle. This static force is given by
@@
where $A_{i}$ and $Z_{i}$ are the atomic mass number and the atomic number of the $i$th particle, respectively, and $\hat{\mathbf{r}}$ is the unit vector pointing from one particle to the other. In an atom interferometer, both of these forces can be detected with a differential acceleration measurement between atomic species (Fig. 3). Detection of the static force sourced by the Earth requires an equivalence principle test in which systematic errors are controlled, while the time-oscillating force can in principle be detected even without accounting for DC systematic effects.
@@
Figure 6: Constraint plot and projected sensitivities to vector dark matter coupled to $B-L$. JHU (DC): projected sensitivity of static equivalence principle test in JHU apparatus with interferometer geometry shown in Fig. 3. JHU (AC): projected sensitivity of search for time-dependent differential acceleration in JHU apparatus. MAGIS-100 (DC): projected sensitivity of static equivalence principle test with Yb isotopes in MAGIS-100 long-baseline interferometer. MAGIS-100 (AC): projected sensitivity of search for time-dependent differential acceleration with Yb isotopes in MAGIS-100. Also shown are existing constraints from the MICROSCOPE experiment Touboul et al. (2022) (purple curve), the Eöt-Wash experiment Shaw et al. (2022) (green curve), and the LIGO and VIRGO gravitational-wave detectors Abbott (2022). Shaded regions are excluded by existing constraints. All curves are 95% confidence limits.

### v3

@@ Abstract
We analyze the projected sensitivity of a laboratory-scale ytterbium atom interferometer to scalar, vector, and pseudoscalar dark matter signals. A frequency ratio measurement between two transitions in 171Yb enables a search for variations of the fine-structure constant that could surpass existing limits by a factor of 100 in the mass range $10^{-22}$ eV to $10^{-16}$ eV. Differential accelerometry between ytterbium isotopes yields projected sensitivities to scalar and vector dark matter couplings that are stronger than the limits set by the MICROSCOPE equivalence principle test, and an analogous measurement in the MAGIS-100 long-baseline interferometer would be more sensitive than previous bounds by factors of 10 or more. A search for anomalous spin torque in MAGIS-100 is projected to reach similar sensitivity to atomic magnetometry experiments. We discuss strategies for mitigating the main systematic effects in each measurement. These results indicate that improved dark matter searches with ytterbium atom interferometry are technically feasible.
@@
Phenomenologically, the signal in a dark matter direct detection experiment changes qualitatively at a dark matter particle mass of $\sim 1$ eV. If the dark matter particle mass is above $1$ eV, the phase space density of the dark matter near the Earth is less than one, and the dark matter is expected to exhibit particle-like behavior (e.g., scattering off of ordinary matter). Several experiments Alkhatib et al. (2021); Abdallah et al. (2015) have searched for dark matter in this regime. If the dark matter particle mass is below $1$ eV, however, the phase space density of the dark matter near the Earth is greater than one. In this case, the dark matter must be a boson, and its behavior is analogous to that of a classical field: all of its lowest-order couplings to the Standard Model yield signals that oscillate at the Compton frequency of the dark matter Graham et al. (2016a); Arvanitaki et al. (2015). Experiments searching for this “ultralight” dark matter include optical clocks Filzinger et al. (2023); Sherrill et al. (2023), microwave cavity experiments Bartram et al. (2021), nuclear magnetic resonance experiments Aybas et al. (2021); Wu et al. (2019), and atomic magnetometers Bloch et al. (2020); Lee et al. (2023), among others. Ultralight dark matter would also give rise to new static forces between ordinary matter footnote1 . Such forces have been constrained by equivalence principle tests Touboul et al. (2022); Schlamminger et al. (2008).
@@
In this work, we consider the potential physics reach of dark matter detection experiments based on atom interferometry. In a light-pulse atom interferometer Hogan et al. (2009), ultracold atoms are split by atom-light interactions into a superposition of external states, which are then recombined and interfered. Depending on the interferometer geometry, the phase of an atom interferometer can be sensitive to inertial forces Kasevich and Chu (1991), recoil velocity Bordé (1989), and/or the energy difference between internal states Dimopoulos2008 . Atom interferometers have been used to test the equivalence principle at a relative accuracy of about $10^{-12}$ Asenbaum et al. (2020), measure the fine-structure constant Morel et al. (2020); Parker et al. (2018), and observe gravitational phase shifts in nonlocal quantum systems Overstreet et al. (2022); Asenbaum et al. (2017). In addition, several gravitational wave detectors based on atom interferometry Yijun Jiang et al. (2021); Badurina et al. (2020); Canuel et al. (2018) are under construction.
@@
High-precision atom interferometers typically have a measurement time on the order of $1$ s and a cycle time of a few seconds; thus, atom interferometers are naturally sensitive to dark matter with Compton frequency $\lesssim 1$ Hz. With a measurement campaign of about one year, an atom interferometry experiment can search for oscillating dark matter signals over eight orders of magnitude in the dark matter particle mass (from $10^{-8}$ Hz to $1$ Hz, or $10^{-22}$ eV to $10^{-14}$ eV). The possible signals from ultralight dark matter depend on its spin and parity. A spin-zero or spin-one dark matter field can produce six qualitatively different signals at lowest order Graham et al. (2016a). Of these, atom interferometers can be sensitive to three: variations of fundamental constants, accelerations, and spin torques. In addition, atom interferometers can search for static forces induced by dark matter.
@@
One of the main considerations for an atom-interferometric dark matter detection experiment is the choice of atomic species. Ideally, the experiment should use an atom that provides high sensitivity to dark matter signals while minimizing systematic effects. Here we consider atom interferometry with ytterbium isotopes. Ytterbium offers the highest sensitivity of any neutral atom to variations of the fine-structure constant in one of its excited states Safronova et al. (2018); Dzuba et al. (2018), possesses multiple isotopes that can be cooled simultaneously for a differential acceleration measurement Kitagawa et al. (2008), and has an isotope with nuclear spin $1/2$ that can be used to detect spin torque. In addition, the alkaline-earth-like electronic structure of ytterbium provides magnetic insensitivity in the ${}^{1}S_{0}$ electronic ground state, and its multiple narrow-linewidth transitions facilitate precise measurements of transition frequencies. Atom interferometry with ytterbium has previously been demonstrated Gochnauer et al. (2021) and has previously been proposed for tests of fundamental physics Hartwig et al. (2015).
@@
The remainder of this paper is organized as follows. The apparatus is described in Section II, and its sensitivities to scalar, vector, and pseudoscalar dark matter are discussed in Sections III.A, III.B, and III.C, respectively. Section IV describes the main systematic effects in these measurements and how they will be controlled. Section V compares atom-interferometric dark matter searches to other experiments and discusses the prospects for future sensitivity improvements.
@@
The JHU experimental apparatus will consist of a source of ultracold ytterbium and an atomic fountain in which the atoms can freely fall during interferometry sequences. Clouds of ultracold ytterbium will be produced in a 3D magneto-optical trap (MOT) that is loaded by a commercial Zeeman slower/2D MOT. The 3D MOT will utilize core-shell techniques Lee et al. (2015) to increase atom number and phase space density. The atoms will be evaporatively cooled in an optical dipole trap and then launched into a magnetically shielded atomic fountain by an optical lattice. The height of the magnetically shielded region will be $2.5$ m, enabling interferometry over a free-fall distance of $2$ m. After the interferometry sequence is complete, the interferometer phase will be measured via fluorescence detection of the number of atoms in each output port.
@@
For the purpose of generating estimated sensitivities to dark matter, we assume that $10^{5}$ atoms will participate in each measurement. We also assume that the coherence time of each measurement will be $1.3$ s, limited by the available free fall time, and that the experimental cycle time will be $10$ s. These atom numbers and cycle times have previously been achieved in high-precision atom interferometers Asenbaum et al. (2020). Finally, we assume that each measurement campaign will have a duration of one year.
@@
Atom interferometers naturally detect energy level shifts and accelerations. With the appropriate geometry, an atom interferometer can be sensitive to scalar, vector, or pseudoscalar (axion-like) couplings to ordinary matter. In this section, we describe interferometer geometries suitable for detecting each of these dark matter candidates and calculate the projected sensitivity of the JHU apparatus to the associated coupling. For each estimate, we assume shot-noise-limited statistics and a one-year measurement campaign. The leading systematic effects for each measurement are discussed in Section IV.
@@
$\alpha(t)=\alpha_{0}\left(1+d_{e}\varphi_{0}\cos\left(\frac{m_{\text{DM}}c^{2}}{\hbar}t\right)\right)$ (3)
@@
Note that $\Delta q$ describes the relativistic corrections to the transition frequency Dzuba et al. (2018). A frequency ratio measurement between two transitions is sensitive to variations of $\alpha$ as long as the value of $\Delta q$ differs between them, and transitions with large values of $\Delta q$ have the highest discovery potential.
@@
We propose to use a 171Yb atom interferometer to measure the frequency ratio between the ${}^{1}S_{0}\leftrightarrow{}^{3}P_{0}$ transition at $578$ nm and the ${}^{1}S_{0}\leftrightarrow 4\text{f}^{13}6\text{s}^{2}5\text{d}\,(J=2)$ transition at $431$ nm. The energy of the $4\text{f}^{13}6\text{s}^{2}5\text{d}\,(J=2)$ state has the highest sensitivity among neutral-atom states to variations of $\alpha$ Safronova et al. (2018). An interferometer geometry for this frequency comparison is depicted in Fig. 1. In each of the four Ramsey-Bordé interferometers, a single-photon transition creates a superposition of internal states, which are interfered to produce an interferometer phase of the form Hogan et al. (2009); Dimopoulos2008
$\phi=2(\omega-\omega_{L})T+\frac{\hbar\omega_{L}}{mc^{2}}(2\omega-\omega_{L})T+\frac{\omega}{c}gT^{2}+\cdots.$ (5)
Here $\omega$ is the transition frequency, $\omega_{L}$ is the laser frequency, $T$ is the interferometer time, $m$ is the atomic mass, and $g$ is the magnitude of the laboratory acceleration relative to a freely falling geodesic. The first term is the desired signal, while the second and third terms represent systematic effects due to the recoil velocity and the laboratory acceleration, respectively footnote3 . To suppress these effects, we consider the differential phase between pairs of interferometers [(a) - (b) and (c) - (d)], which has the form
[…]
and is insensitive to the phase shifts arising from the recoil velocity and the laboratory acceleration because they are common to both interferometers. In Eq. 6, we have explicitly included the time dependence of $\omega$. For most of the dark matter parameter space we consider, the oscillation period of $\omega$ is much longer than the interferometer time, and $\omega$ is approximately constant during each measurement. For higher dark matter masses, the oscillation of $\omega$ during a single measurement becomes significant and leads to loss of sensitivity. In addition, we note that this phase shift has the same form as that of an optical lattice clock interrogated with a Ramsey sequence. As in an optical clock, the population difference between the two output states will be a function of the frequency difference between the atomic transition and the laser.
Both of the single-photon transitions depicted in Fig. 1 can be driven with technically feasible laser systems. For the ${}^{1}S_{0}\leftrightarrow{}^{3}P_{0}$ transition at $578$ nm, which is weakly allowed in 171Yb due to the nonzero nuclear spin, a laser intensity of $1$ W/cm2 is sufficient to obtain a Rabi frequency of $40$ kHz. The ${}^{1}S_{0}\leftrightarrow 4\text{f}^{13}6\text{s}^{2}5\text{d}\,(J=2)$ transition at $431$ nm is an M2 transition with Rabi frequency $17$ Hz $\cdot\,\sqrt{I}$/(mW/cm2)1/2 Dzuba et al. (2018). Thus, a Watt-class laser with roughly $1$ cm beam size would enable $1$ kHz Rabi frequency. The lifetime of the $4\text{f}^{13}6\text{s}^{2}5\text{d}\,(J=2)$ state has been estimated to be between $60$ s and $200$ s, limited by decay to the triplet $P$ manifold Safronova et al. (2018); Dzuba et al. (2018).
An alternative geometry for measuring this frequency ratio is shown in Fig. 2, where each beamsplitter is implemented by a Doppler-free two-photon transition driven by counterpropagating laser beams Alden (2014). Compared to the geometry in Fig. 1, Doppler-free transitions naturally suppress the recoil shift, thereby reducing the number of interferometers required from four to two, and prevent the atom loss into undesired momentum states that is intrinsic to Ramsey-Bordé interferometers. However, Doppler-free two-photon transitions require higher laser power than single-photon transitions for a given Rabi frequency.
@@
Here $r$ is the distance between particles, $m_{i}$ and $Q_{i}$ are the mass and the dilaton charge of particle $i$, respectively, and $\lambda$ is the Compton wavelength of the dark matter particle. Note that the coupling constant $d_{e}$ in this expression is the same quantity as in Eq. 1. Since the dilaton charge of an atom is a function of its mass number and atomic number (see Appendix A.I), this potential leads to a composition-dependent static force
@@
where $n$ is the number of photon recoils in the initial beamsplitter, $k$ is the magnitude of the laser wavevector, and $\Delta g$ is the acceleration difference between isotopes projected onto the interferometer axis. For this measurement, we consider a comparison between 170Yb and 176Yb, which are both bosons and have favorable scattering lengths Kitagawa et al. (2008) that allow simultaneous evaporative cooling. The beamsplitters can be implemented by Bragg transitions on the electric dipole transition at $399$ nm or Bragg transitions on the intercombination transition at $556$ nm. Either sequential two-photon transitions Asenbaum et al. (2020) or multi-photon transitions Parker et al. (2018) could be used to impart momentum to the atoms. We assume $n=500$, allowing a maximum wave packet separation of $2$ m at time $T$. Beamsplitters with hundreds of photon recoils have previously been demonstrated Wilkason et al. (2022); Béguin et al. (2023). With these parameters, the JHU apparatus is projected to reach a sensitivity to the Eötvös parameter $\eta\equiv\Delta g/g=2\times 10^{-16}$ after a one-year measurement campaign.
@@
Figure 4: Constraint plot and projected sensitivities to scalar dark matter coupled to the electromagnetic field tensor. JHU $\alpha(t)$: projected sensitivity of spectroscopic search in JHU apparatus with interferometer geometry shown in Fig. 1. JHU (DC): projected sensitivity of static equivalence principle test in JHU apparatus with interferometer geometry shown in Fig. 3. MAGIS-100 (DC): projected sensitivity of static equivalence principle test with Yb isotopes in MAGIS-100 long-baseline interferometer. MAGIS-100 Gradiometer: projected sensitivity of gradiometric measurement proposed in Ref. Yijun Jiang et al. (2021), assuming the same phase resolution and interferometer parameters as other projections. Also shown are existing constraints from the MICROSCOPE experiment Touboul et al. (2022) (purple curve) and Yb ion clock experiments at NPL Sherrill et al. (2023) (green curve) and PTB Filzinger et al. (2023) (yellow curve). Shaded regions are excluded by existing constraints. All curves are 95% confidence limits.
@@
Figure 5: Constraint plot and projected sensitivities to scalar dark matter coupled to the electron mass. JHU (DC): projected sensitivity of static equivalence principle test in JHU apparatus with interferometer geometry shown in Fig. 3. MAGIS-100 (DC): projected sensitivity of static equivalence principle test with Yb isotopes in MAGIS-100 long-baseline interferometer. MAGIS-100 Gradiometer: projected sensitivity of gradiometric measurement proposed in Ref. Yijun Jiang et al. (2021), assuming the same phase resolution and interferometer parameters as other projections. Also shown are existing constraints from the MICROSCOPE experiment Touboul et al. (2022) (purple curve) and a hydrogen maser-optical cavity comparison Kennedy et al. (2020) (green curve). Shaded regions are excluded by existing constraints. All curves are 95% confidence limits.
@@
The projected sensitivities of these measurements to the scalar dark matter coupling constant $d_{e}$ are shown in Fig. 4 as a function of the dark matter particle mass. The projections are compared to existing constraints from ytterbium ion clock experiments Filzinger et al. (2023); Sherrill et al. (2023) and the MICROSCOPE space-based equivalence principle test Touboul et al. (2022). We calculate the projected sensitivity to variations of $\alpha$ both analytically and by taking the discrete Fourier transform of simulated data. In the JHU experiment, the Lomb-Scargle periodogram VanderPlas (2018) will be used instead of the discrete Fourier transform to optimally account for variations in cycle time and experimental dead time. The projections take into account the stochastic amplitude of the dark matter field, which diminishes sensitivity to oscillating signals by a factor of about 3 when the measurement campaign is shorter than the dark matter coherence time Centers et al. (2021). We conservatively apply this sensitivity reduction across the entire parameter space. A more sophisticated analysis, following Ref. Centers et al. (2021), would improve the projected sensitivity at frequencies above $10^{-1}$ Hz by a factor of order one.
At low dark matter masses, the projected sensitivity of the search for variations of $\alpha$ scales like $1/m_{\text{DM}}$ (see Eqs. 2 and 3), while the projected sensitivity of the equivalence principle test becomes independent of $m_{\text{DM}}$ (see Eq. 8). At higher dark matter masses, the search for variations of $\alpha$ loses sensitivity due to signal averaging over the $\sim 1$ s coherence time of each experimental run. To avoid the loss of sensitivity when $m_{\text{DM}}c^{2}/\hbar\cdot T=n\pi$ for integer $n$, the interferometer time $T$ can be varied, as described in Ref. Badurina (2022). When $m_{\text{DM}}c^{2}/\hbar$ exceeds the Nyquist frequency, searches for dark matter can be performed using the techniques introduced in Ref. Badurina (2023), which we do not consider here. In contrast, the equivalence principle test loses sensitivity at higher dark matter masses due to the exponential decay term in Eq. 8, which effectively limits the quantity of material in the Earth that sources the static force. The change in slope at $m_{\text{DM}}\sim 10^{-13}$ eV arises from our model of the Earth, which takes into account the composition difference between the core and the mantle (see Appendix A.II).
We also plot the projected sensitivity of an equivalence principle test performed between 170Yb and 176Yb in the MAGIS-100 long-baseline interferometer at Fermilab Yijun Jiang et al. (2021). For this projection, we use the interferometer geometry in Fig. 3 and the same experimental parameters as in the JHU apparatus (atom number, photon recoils, etc.), except that the interferometer time is increased from $T=0.64$ s to $T=4.5$ s. We note that this measurement in MAGIS-100 would reach an Eötvös parameter sensitivity of $\eta=5\times 10^{-18}$.
@@
The JHU atom-interferometric search for variations of the fine-structure constant is projected to improve on existing experiments by two orders of magnitude in the mass range $10^{-22}$ eV to $10^{-16}$ eV. The ytterbium equivalence principle test performed in the JHU apparatus is projected to reach a similar sensitivity to the MICROSCOPE experiment, while the MAGIS-100 equivalence principle test would be a factor of 10 more sensitive than existing bounds from $10^{-17}$ eV to $10^{-12}$ eV.
@@
We note that a gradiometer utilizing a single optical transition has also been proposed to search for scalar dark matter Arvanitaki et al. (2018); Yijun Jiang et al. (2021). The projected sensitivity of such an experiment in MAGIS-100 to $d_{e}$ and $d_{m_{e}}$ is shown in Figs. 4 and 5, respectively. A ytterbium gradiometer would have similar projected sensitivity to a strontium gradiometer for these searches. Nevertheless, we project that the gradiometric operating mode would be less sensitive than the bounds set by the MICROSCOPE experiment throughout the mass range. The discrepancy between our projection and previous work Arvanitaki et al. (2018); Yijun Jiang et al. (2021) is explained by our more conservative estimate of the attainable phase resolution footnote2 . In addition, we note that gradiometric searches for dark matter at low frequencies are severely constrained by gravity gradient noise, which becomes challenging to characterize at frequencies below $1$ Hz Badurina (2023). Since the differential phase response to gravity gradients scales linearly with the spatial separation between two interferometers, the co-located interferometers in this proposal are much less susceptible to gravity gradient noise.
@@
where $\hat{\mathbf{e}}$ points in the polarization direction of the dark matter field. In the presence of a time-varying acceleration $g(t)$, the phase response of a Mach-Zehnder interferometer is given by Overstreet (2021)
[…]
In addition, atoms can exert a force on one another by exchanging the vector particle. This static force is given by
@@
where $A_{i}$ and $Z_{i}$ are the mass number and the atomic number of the $i$th particle, respectively, and $\hat{\mathbf{r}}$ is the unit vector pointing from one particle to the other. In an atom interferometer, both of these forces can be detected with a differential acceleration measurement between atomic species (Fig. 3). Detection of the static force sourced by the Earth requires an equivalence principle test in which systematic errors are controlled, while the time-oscillating force can in principle be detected even without accounting for DC systematic effects.
@@
Figure 6: Constraint plot and projected sensitivities to vector dark matter coupled to $B-L$. JHU (DC): projected sensitivity of static equivalence principle test in JHU apparatus with interferometer geometry shown in Fig. 3. JHU (AC): projected sensitivity of search for time-dependent differential acceleration in JHU apparatus. MAGIS-100 (DC): projected sensitivity of static equivalence principle test with Yb isotopes in MAGIS-100 long-baseline interferometer. MAGIS-100 (AC, Yb) and (AC, Sr): projected sensitivities of searches for time-dependent differential acceleration in MAGIS-100 with Yb isotopes and Sr isotopes Yijun Jiang et al. (2021), respectively, assuming the same phase resolution and interferometer parameters as other projections. Also shown are existing constraints from the MICROSCOPE experiment Touboul et al. (2022) (purple curve), the Eöt-Wash experiment Shaw et al. (2022) (green curve), and the LIGO and VIRGO gravitational-wave detectors Abbott (2022). Shaded regions are excluded by existing constraints. All curves are 95% confidence limits.